Upton and Verlander Company uses a job-order costing system. Upton and Verlander has two manufacturing departments—Molding and Fabrication. Upton and Verlander provided the following estimates at the beginning of the year:

  

Molding Fabrication Total
Machine-hours 30,000 40,000 70,000
Fixed manufacturing overhead costs $ 790,000 $ 240,000 $ 1,030,000
Variable manufacturing overhead cost per machine-hour $ 5.30 $ 5.30

  

During the year, the company had no beginning or ending inventories and it started, completed, and sold only two jobs—Job D-70 and Job C-200. It provided the following information related to those two jobs:

  

Job D-70: Molding Fabrication Total
Direct materials cost $ 376,000 $ 327,000 $ 703,000
Direct labor cost $ 220,000 $ 140,000 $ 360,000
Machine-hours 23,000 7,000 30,000

  

Job C-200: Molding Fabrication Total
Direct materials cost $ 200,000 $ 280,000 $ 480,000
Direct labor cost $ 140,000 $ 240,000 $ 380,000
Machine-hours 7,000 33,000 40,000

Upton and Verlander had no underapplied or overapplied manufacturing overhead during the year.

Required:

1. Assume Upton and Verlander uses a plantwide predetermined overhead rate based on machine-hours.

a. Compute the plantwide predetermined overhead rate.

b. Compute the total manufacturing cost assigned to Job D-70 and Job C-200.

c. If Upton and Verlander establishes bid prices that are 140% of total manufacturing cost, what bid prices would it have established for Job D-70 and Job C-200?

d. What is Upton and Verlander’s cost of goods sold for the year?
Answer #1
1a
Estimated Variable manufacturing overhead 371000 =(30000*5.3)+(40000*5.3)
Estimated Fixed manufacturing overhead 1030000
Total Estimated manufacturing overhead 1401000
Divide by Estimated Machine hours 70000
Predetermined overhead rate 20.01 per MH
1b
Job D-70 Job C-200
Direct materials cost 703000 480000
Direct labor cost 360000 380000
Overhead applied 600300 800400
Total manufacturing cost 1663300 1660400
1c
Job D-70 Job C-200
Total manufacturing cost 1663300 1660400
Add: Markup @40% 665320 664160
Bid prices 2328620 2324560
1d
Total manufacturing cost Job D-70 1663300
Total manufacturing cost Job C-200 1660400
Cost of goods sold 3323700
Workings:
Job D-70 Job C-200
Overhead applied =30000*20.01 =40000*20.01